use susstitution to solve each system of equations.
0.5x-2y=17 2x+y=104

Answer by whty89(20)

EQUATION#1.).5x-2y=17 EQUATION#2.)2x+y=104
1.)solve for y in Eq#2 by subtracting 2x from both sides. y=-2x+104
2.)substitute -2x+104 for y into EQ#1.
3.).5x-2(-2x+104)=17
.5x+4x-208=17
4.5x-208=17
4.5x=225
x=50
4.)substitute 50 for x into y=-2x+104
y=-2(50)+104
y=-100+104
y=4
5.) the solution set is (50,4)